What is electric flux,electric field and electric potential inside the charged sphere?


1
Expert's answer
2021-06-07T19:22:55-0400

Electric Field:

According to Gauss’s law, the electric field inside a conducting sphere is zero. This is due to the fact that the total charge enclosed is zero.


Electric Flux:

Electric Flux in a sphere is given by:


"Electric Flux = E * A = E * 4 * \pi * r^{2}"


As the electric field inside a sphere is zero, Therefore

Electric Flux = E * A = 0


Therefore, electric flux is zero inside a sphere.


Electric Potential:

The electric potential is constant inside a sphere. And it is same as that on the surface of conductor.